### v2

# SEA: Spectral Edge Attack

Abstract
Graph-based machine learning algorithms occupy an important position in today’s AI landscape. The ability of graph topology to represent complex data structures is both the key strength of graph algorithms and a source of their vulnerability. In other words, attacking or perturbing a graph can severely degrade the performance of graph-based methods. For the attack methods, the greatest challenge is achieving strong attack effectiveness while remaining undetected. To address this problem, this paper proposes a new attack model that employs spectral adversarial robustness evaluation to quantitatively analyze the vulnerability of each edge in a graph under attack. By precisely targeting the weakest links, the proposed approach achieves the maximum attack impact with minimal perturbation. Experimental results demonstrate the effectiveness of the proposed method.
footnotetext: Correspondence to: yongyuw@mtu.edu
1 Introduction
Many important datasets are naturally represented as graphs, spanning online social and recommendation platforms (Facebook, Amazon), molecular interaction resources such as BioGRID, and citation or document-link networks like PubMed and Arxiv. Given the prevalence of such data, a number of widely used tasks and algorithms have been developed accordingly. From the task perspective, the most common problems include classification and clustering. From the algorithmic or modeling perspective, representative approaches include graph neural networks and spectral clustering.
The broad deployment of these tasks and algorithms also makes their security a critical concern. As stated in the ancient Chinese military treatise The Art of War, “know yourself and know your enemy, and you will never be defeated.” Without strong, well-designed attack methodologies, it is difficult to meaningfully evaluate defensive techniques or drive their improvement.
Motivated by this perspective, this paper proposes a new high-strength attack method. The key idea is to perform spectral adversarial robustness evaluation on the graph structure before launching the attack, thereby quantitatively estimating the vulnerability of each edge. This enables the attacker to pinpoint and target the weakest links, achieving maximal impact with only a very small number of perturbations to undermine the system while remaining difficult to detect.
2 Preliminaries
Spectral graph theory analyzes the structure of a graph through the eigenvalues and eigenvectors of matrices derived from it, most notably the graph Laplacian. Let $G=(V,E,w)$ be a weighted graph, where $V$ and $E$ denote the vertex and edge sets, respectively, and $w:E\to\mathbb{R}_{>0}$ assigns a positive weight to each edge. The (combinatorial) Laplacian $L$ of $G$ is a symmetric, diagonally dominant (SDD) matrix defined entry-wise by
$L(p,q)=\begin{cases}-w(p,q)&\text{if }(p,q)\in E,\\ \sum\limits_{(p,t)\in E}w(p,t)&\text{if }p=q,\\ 0&\text{otherwise}.\end{cases}$ (1)
To improve numerical stability and to factor out global scaling effects, it is common to work with the normalized Laplacian
$L_{\text{norm}}=D^{-1/2}LD^{-1/2},$
where $D$ is the diagonal matrix of (weighted) vertex degrees.
Many spectral methods require only a small portion of the spectrum rather than a full eigendecomposition [2]. For example, in clustering-oriented applications, the smallest Laplacian eigenvalues and their corresponding eigenvectors typically capture the most salient community information [3]. As a result, computing all eigenpairs is unnecessary and becomes infeasible for large graphs. A standard tool for extracting extremal eigenvalues is the Courant–Fischer minimax principle [4], which gives a variational characterization suitable for iterative algorithms. Concretely, for a Laplacian $L\in\mathbb{R}^{|V|\times|V|}$, the $k$-th eigenvalue can be written as
$\lambda_{k}(L)=\min_{\dim(U)=k}\ \max_{x\in U,\ x\neq 0}\ \frac{x^{\top}Lx}{x^{\top}x}.$
In a number of settings, one is interested in relating two matrices rather than studying a single Laplacian in isolation. The Courant–Fischer characterization extends naturally to a generalized form. Let $L_{X},L_{Y}\in\mathbb{R}^{|V|\times|V|}$ be two Laplacian matrices satisfying $\mathrm{null}(L_{Y})\subseteq\mathrm{null}(L_{X})$. Then, for $1\leq k\leq\mathrm{rank}(L_{Y})$, the $k$-th eigenvalue of $L_{Y}^{+}L_{X}$ admits the variational representation
$\lambda_{k}(L_{Y}^{+}L_{X})=\min_{\begin{subarray}{c}\dim(U)=k,\\ U\perp\mathrm{null}(L_{Y})\end{subarray}}\ \max_{x\in U,\ x\neq 0}\ \frac{x^{\top}L_{X}x}{x^{\top}L_{Y}x},$
where $L_{Y}^{+}$ denotes the Moore–Penrose pseudoinverse of $L_{Y}$.
3 Method
3.1 Phase 1: Learning Latent Representations
Machine learning models can be viewed as transformation engines that convert inputs into outputs by successively reshaping representations. For example, deep neural networks compress and refine raw feature vectors through multiple nonlinear layers. Likewise, methods such as Support Vector Machines (SVM) and Support Vector Clustering (SVC) employ kernel functions to implicitly lift data from the original feature space into a higher-dimensional one [1]. Adversarial attacks seek to exploit this mechanism by inducing failures in the input–output transformation, causing the model to produce incorrect predictions. Therefore, we first run the model in its standard (benign) setting to obtain the model-induced transformation of the input feature space, i.e., the resulting latent representations.
3.2 Phase 2: Capturing the Underlying Manifold of the Latent Representation
To analyze the latent representations obtained in Phase 1 through spectral graph theory, we need to build a graph that captures the underlying manifold of these embeddings. Among the many graph construction strategies, the $k$-nearest neighbor ($k$-NN) graph is one of the most widely used and is well suited for manifold approximation. Accordingly, in this phase we construct a $k$-NN graph from the transformed feature matrix produced by the trained model.
3.3 Phase 3: Adversarial Robustness Evaluation
Let $L_{X}$ be the Laplacian of the original $k$-NN graph and $L_{Y}$ the Laplacian of the $k$-NN graph built in the spectrally embedded space. The model-induced sensitivity can be measured by examining the leading generalized eigenstructure of $L_{Y}^{+}L_{X}$. In particular, we construct a weighted eigen-subspace $V_{s}\in\mathbb{R}^{|V|\times s}$ that yields a spectral embedding of the input-manifold graph $G_{X}=(V,E_{X})$, where $|V|$ is the number of nodes:
$V_{s}=\begin{bmatrix}v_{1}\sqrt{\zeta_{1}},\;v_{2}\sqrt{\zeta_{2}},\;\ldots,\;v_{s}\sqrt{\zeta_{s}}\end{bmatrix}$.
Here, $\zeta_{1}\geq\zeta_{2}\geq\cdots\geq\zeta_{s}$ are the top $s$ eigenvalues of $L_{Y}^{+}L_{X}$, and $v_{1},\ldots,v_{s}$ are their associated eigenvectors.
The embedding of node $p\in V$ is given by the $p$-th row of $V_{s}$, resulting in an $s$-dimensional representation. Following [5], we quantify the stability of an edge $(p,q)\in E_{X}$ using the spectral distance induced by this embedding:
$\bigl\|V_{s}^{\top}e_{p,q}\bigr\|_{2}^{2}.$
Motivated by this measure, we define the Spade score of edge $(p,q)$ as
$\mathrm{Spade}(p,q)=\bigl\|V_{s}^{\top}e_{p,q}\bigr\|_{2}^{2},$
where larger values indicate lower robustness, i.e., edges that are more vulnerable to perturbations aligned with the directions implied by their incident nodes.
3.4 Phase 4: Identifying Weak Links for Attack
For the edges in the graph used by the machine learning model, we first rank them in descending order by their Spade scores and select the top fraction to form a non-robust edge set. Since these edges are most susceptible to perturbations, we focus our attack on this set.
4 Experiments
In this section, we use the method proposed in this paper to attack two types of dominant models in the field of graph-based machine learning: graph neural networks and spectral clustering. For the selected edges to be attacked, our algorithm randomly assigns their weights to an integer value within a specified range.To demonstrate the effectiveness of our method, we use random sampling to select the same number of edges from the graph as an attack baseline. We then perturb their weights by assigning random values drawn from the same range used in our method, and compare the resulting attack impact against ours.
4.1 Attacking GNNs
The results are shown in Table 1
Table 1: Baseline test accuracy and test accuracy drops under edge-weight randomization attacks (Spade-selected vs. Random-selected). Random results are averaged over 20 runs (mean $\pm$ std).
Dataset Model Baseline (Test) Drop (Ours) Drop (Random, mean $\pm$ std) Ours (wins/20)
Cora GAT 0.8826 0.0367 0.0138 $\pm$ 0.0083 20/20
Cora GraphSAGE 0.8936 0.0202 0.0114 $\pm$ 0.0074 18/20
PubMed GAT 0.8058 0.0124 0.0090 $\pm$ 0.0032 18/20
PubMed GraphSAGE 0.8489 0.0061 0.0038 $\pm$ 0.0018 18/20
4.2 Attacking Spectral Clustering
The results are shown in Table 2
Table 2: USPS data set: ACC drops (percentage points). Larger drop indicates a stronger attack.
change_ratio Baseline ACC Our Drop Random Drop (mean $\pm$ std) Ours (wins/20)
0.01 66.4229 +3.5492 $1.9542\pm 3.8065$ 20/20
0.02 66.4121 +2.1402 $1.2363\pm 3.5204$ 20/20
0.03 66.4229 +1.7853 $1.4390\pm 2.9963$ 20/20
0.04 66.4229 +3.4846 $0.6458\pm 0.4842$ 20/20
0.05 66.4121 +3.3125 $1.2266\pm 2.8007$ 20/20
5 Conclusion
This paper proposes a novel method for precisely attacking the graphs used in graph-based machine learning algorithms, based on spectral adversarial robustness evaluation. Experiments on representative graph algorithms from neural networks and statistical learning—namely graph neural networks (GNNs) and spectral clustering—demonstrate that the proposed method is highly effective.